20Cr = 20C(r+2), find the value of r.

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

20Cr = 20Cr +2, find the value of r.
----
20Cr = 20Cr +2
0 = 2
No solution.

Answer by Edwin McCravy(20056)


For all positive integers n and r, where n ≥ r ≥ 0
nCk = nC(n-k)

If you mean 20Cr = 20C(r+2), then since r+2 > r, we turn nCk = nC(n-k)
around as
  
 nC(n-k) =  nCk
20Cr     = 20C(r+2)

n=20, r=n-k, k=r+2
      r=20-k
      r=20-(r+2)
      r=20-r-2
      r=18-r
     2r=18
      r=9

Checking
  20Cr = 20C(r+2)
  20C9 = 20C(9+2)
  20C9 = 20C11
167960 = 167960

Edwin